OpenStudy (anonymous):

Factor expression below. Write as a polynomial in descending order. 16x^2-1

OpenStudy (whpalmer4):

That's a difference of two squares: \[(a-b)(a+b) = a^2-ba+ba-b^2 = a^2-b^2\]

OpenStudy (anonymous):

still confused....

OpenStudy (whpalmer4):

Well, what is the square root of 1?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Good. What is the square root of \(16x^2\)?

OpenStudy (anonymous):

4x??

OpenStudy (whpalmer4):

Yes! So now do you see why I called it a difference of two squares?

OpenStudy (whpalmer4):

It's \((4x)^2 - (1)^2\), right?

OpenStudy (anonymous):

could u help me with 9 more questions?? I need them to graduate in 24 hours...

OpenStudy (whpalmer4):

If you have a difference of two squares, \(a^2-b^2\), you can factor that as \[(a+b)(a-b)\] Here we have \(a=4x\) and \(b = 1\) so it factors as \[(16x^2-1) = (4x-1)(4x+1)\]

OpenStudy (whpalmer4):

What's the next question?

OpenStudy (anonymous):

left side of the attachment

OpenStudy (whpalmer4):

Here we've got a difference of squares going in the other direction: we've got the factors, and we'll get the difference of squares. \[\frac{x+3}{x+2}*\frac{x-3}{x-2}\]What does the numerator become after multiplying? How about the denominator? Remember, \((a+b)(a-b) = a^2-b^2\)

OpenStudy (anonymous):

sooo B??

OpenStudy (whpalmer4):

yes!

OpenStudy (whpalmer4):

1 more, then I have to go...

OpenStudy (anonymous):

okay sounds good left side again

OpenStudy (whpalmer4):

A couple of ways you could go here. Just read intersection point off the graph, or set the two equations equal to each other and solve for \(x\). \[\frac{3}{5}x-1 = x-3\]Multiply through by 5\[3x-5=5x-15\]Subtract 3x from both sides, add 15 to both sides:\[10=2x\]\[x=5\]now plug that into either formula to get \(y\)\[y=x-3=5-3=2\]so our point of intersection is \((5,2)\) and if you look at the graph, I think you'll agree that is where the lines intersect.

OpenStudy (anonymous):

thank you so much u have been a great help!

OpenStudy (whpalmer4):

Hopefully if you've got any more of those types of problems you won't have any trouble following the example...good luck!